# Pushing the limits of eccentricity in planet-disc interactions

### v1

@@ Abstract
Modelling the gravitational interaction between an eccentric perturber and a differentially shearing gas disc is a longstanding problem with various astrophysical applications, ranging from the evolution of planetary systems to the migration of black holes embedded in AGN discs. Recent work has promoted the use of a global, linear, modal approach for calculating the excited wake and the resulting feedback on the perturber’s orbital evolution. In this work we perform a complementary suite of targeted hydrodynamic simulations to test this linear framework across a range of disc temperature and density profiles. In particular, we push from circular orbits to highly eccentric trajectories for which the perturber moves supersonically with respect to the background gas. We find remarkable agreement between our simulations and the linear methodology across a range of diagnostics – lending support to the predicted wake morphologies, complex radial torque density profiles, and torque reversal phenomena, which occur when the eccentricity exceeds the local scale height. In contrast, comparison with previous fitting functions exposes noticeable discrepancies, cautioning against their indiscriminate use in studies which explore a wide range of perturber eccentricities, in varied disc structures. Our simulations also probe the fundamentally nonlinear effects of shock induced angular momentum deposition and coorbital horseshoe drags, which exhibit clear variations with eccentricity. Finally, this careful comparison between linear theory and numerics provides a detailed benchmark for planet-disc interaction problems and therefore we have provided a repository of our linear calculations for use as a rigorous test of future numerical investigations.
@@ 1 Introduction
Whilst the majority of attention has focused on the regime where the perturber is on a circular orbit, there is evidence to suggest that eccentric orbits might also be common. Indeed, observed spiral arm pitch angles in protoplanetary discs are sometimes difficult to explain using a single, circular planet (e.g. Monnier et al., 2019; Uyama et al., 2020). Furthermore, the presence of multiple spiral pattern speeds in SAO 206462 is suggestive of two planets or an eccentric perturber (Xie et al., 2021). These early planet-disc interactions might also connect with the final exoplanet demographics, which exhibit a clear eccentric population (e.g. Kane et al., 2012; Xie et al., 2016; Eylen et al., 2019; Debras et al., 2021). The origin for these eccentricities might owe to a variety of mechanisms as described in the recent review by Paardekooper et al. (2023), including planet-planet scatterings (e.g. Ford & Rasio, 2008; Lega et al., 2013), secular or resonant gravitational interactions with the surrounding disc (e.g. Goldreich & Sari, 2003; Teyssandier & Ogilvie, 2016; Ragusa et al., 2018) and thermal torques (Masset, 2017; Eklund & Masset, 2017).
@@
Understanding how such eccentricities modify the wave launching process, and associated torque feedback on the perturber, is therefore important. In this work we will focus our attention on the low mass planetary regime where we can gain rigorous analytical insight by means of linearizing the fluid equations. The foundational work in this Type-I regime was laid out by Goldreich & Tremaine (1979, 1980). They argued that the wave excitation and coupling to the planetary potential is localised to Lindblad and corotation resonances. The total torque is then found by a summation over these contributions and only depends on the perturbing potential and disc properties at these discrete locations. Subsequent generalizations refined this picture by accounting for gradients in disc properties (Ward, 1986) and by reintroducing non-WKB terms (Artymowicz, 1993). Nonetheless these methods are fundamentally local in character. An alternative approach is to solve the full, linearized wave equation across the global disc without further approximation, as described by Korycansky & Pollack (1993). This method has proven successful in previous circular planet-disc studies, highlighting appreciable differences compared with the isolated resonant picture (see Rafikov & Petrovich, 2012; Petrovich & Rafikov, 2012), whilst also exemplifying subtle features in the global torque density structure (e.g. Miranda & Rafikov, 2019a, b; Cimerman et al., 2024). Furthermore, extended local approaches have investigated three dimensional effects, whilst also accounting for small orbital eccentricities/inclinations, by means of numerical solutions to modified shearing box wave equations (Tanaka et al., 2002; Tanaka & Ward, 2004; Tanaka & Okada, 2024).
Inspired by these successes, this global methodology was recently extended to eccentric perturber orbits by Fairbairn & Rafikov (2022) (hereafter FR22), which found that the linear theory is capable of capturing the detailed, fine structure seen in the complex wake morphologies simulated by Zhu & Zhang (2022). Subsequently this machinery was employed by Fairbairn & Rafikov (2025) (hereafter FR25) which undertook an expansive study of the torque excitation and associated back-reaction onto the planetary orbital elements. They comprehensively explored a suite of disc profiles and eccentricities up to $0.12$, finding a distinct transition in behaviour when the eccentricity exceeds the local disc aspect ratio. They also improved on previous torque and eccentricity evolution prescriptions found in the literature, which are heuristically motivated and often limited in their range of validity (e.g. Papaloizou & Larwood, 2000; Cresswell & Nelson, 2006; Ida et al., 2020).
@@
These tantalising features demand a more targeted study which can be directly compared with the global, linear method of FR25. Therefore, in this paper we will perform a tailored set of 2D simulations which examine planet-disc interaction across a range of disc profiles whilst pushing towards high eccentricity values of 0.7. This significantly extends the previous upper limit of 0.12 explored in FR25 and challenges the capabilities of the analytical framework. With these bespoke simulations we will examine the wake morphology, radial excitation of torque and the net feedback on the planet, thereby verifying the linear theory in unprecedented detail. In section 2 we will review the methodology underpinning the linear results, before we describe the numerical setup in section 3. We will then present our results in section 4 and discuss them in section 5. Finally we will conclude in section 6. Additionally, in appendix A we discuss the convergence properties of our linear framework and numerical simulations, whilst appendix B provides a repository of our data for benchmarking future numerical investigations.
@@ 2.1 Basic approach
We adopt a simple two-dimensional, coplanar and inviscid disc model about a star of mass $M_{\star}$, described in terms of the cylindrical coordinates $(r,\phi)$. The axisymmetric equilibrium disc is specified by the radial profiles of surface density $\Sigma$ and the vertically-integrated pressure $P=c_{\textrm{s}}^{2}\Sigma$, where $c_{\rm s}$ is the isothermal sound speed. The equation for radial force balance then specifies the angular velocity $\Omega$, such that the azimuthal velocity $u_{\phi}=r\Omega$ while the radial velocity $u_{r}=0$.
We now consider the linear perturbing effect of a planet with mass $M_{\rm p}$, which is inserted into the disc. This linear approach formally requires that $M_{\rm p}\ll M_{\rm th}$, where the thermal mass $M_{\rm th}=h_{\rm p}^{3}M_{\star}$ and $h_{\rm p}$ denotes the disc aspect ratio at the planetary semi-major axis location. Specifically we set this low mass planet on a fixed Keplerian orbit with eccentricity $e$, semi-major axis $a_{\rm p}$ and resulting mean motion $n=\sqrt{GM_{\star}/a_{\rm p}^{3}}$. [^2] A schematic of this setup is visualised in Fig.1 of FR22. Throughout this work, we consider the planet-disc interaction in the astrocentric frame centred on the stellar primary and only include the direct perturbing potential,
$\Phi=-\frac{GM_{\text{p}}}{\sqrt{r^{2}+R^{2}-2rR\cos(\phi-\psi)+\epsilon^{2}}}$ (1)
Here, $R(t)$ and $\psi(t)$ denote the time-dependent planetary coordinates whilst $\epsilon$ is a smoothing length, chosen to soften the singularity at the planet location according to a simple Plummer prescription. In this work, we have neglected the effect of the indirect term arising from the non-inertial frame. Whilst including this term can lead to physically interesting and previously unappreciated features in the disc angular momentum flux (Rafikov et al., in prep ), the dominant wake structure and torque feedback on the planet is not strongly affected (see FR25). Since the orbit is non-circular, one cannot remove the fundamental time-dependence of the direct forcing by simply moving into a frame corotating with the planetary mean motion. Instead, the perturbing potential must be expanded in terms of two Fourier numbers – in azimuthal space and orbital time – such that,
$\Phi=\sum_{m=1}^{\infty}\sum_{l=-\infty}^{\infty}\Phi_{ml}(r)\cos\left(m\phi-lnt\right).$ (2)
Here, the cosine dependence expresses the freedom to choose the phase of the orbit such that pericentre passage occurs at $t=0$ and the perturbing components $\Phi_{ml}$ are real. This external potential forces a linear response in the fluid quantities which are similarly assumed to have a Fourier form,
@@
where $\omega_{ml}=(l/m)n$ denotes the modal pattern speed (generally different from the planetary mean motion when $l\neq m$). Upon linearizing the governing continuity and Euler equations, then inserting the modal ansatz, one arrives at
@@
where $\kappa^{2}=(2\Omega/r)d(r^{2}\Omega)/dr$ is the squared epicylic frequency and $\tilde{\omega}=m(\omega_{ml}-\Omega)$ is the Doppler shifted forcing frequency. Also note that here we have dropped the subscripts $ml$ for ease of notation. Equations (4)-(6) are closed by specifying an equation of state (EoS). Throughout this paper we adopt the locally isothermal approximation for which the temperature $T(r)\propto c_{s}^{2}$ is a fixed function of radius and $\delta P=c_{s}^{2}\delta\Sigma$. We choose to focus on this simple thermodynamic model so as to directly interface with the comparable approach adopted in FR22 and FR25. However, it should be remembered that alternative EoS prescriptions can lead to different wave behaviour – as seen for adiabatic and ‘beta-cooling’ discs by Miranda & Rafikov (2020). Finally, combining equations (4)-(6) in favour of the pseudo-enthalpy variable, $\delta h=\delta P/\Sigma$, eventually yields the master linear equation,
$\frac{d^{2}}{dr^{2}}\delta h$ +$+\left\{\frac{d}{dr}\ln\left(\frac{r\Sigma}{D}\right)-\frac{1}{L_{T}}\right\}\frac{d}{dr}\delta h$
@@
$\left.+\frac{1}{L_{T}}\frac{d}{dr}\ln\left(\frac{r\Sigma}{L_{T}D}\right)+\frac{m^{2}}{r^{2}}+\frac{D}{c_{\textrm{s}^{2}}}\right\}\delta h=$
@@
+$+\left\{\frac{2m\Omega}{r\tilde{\omega}}\left[\frac{d}{dr}\ln\left(\frac{\Sigma\Omega}{D}\right)\right]+\frac{m^{2}}{r^{2}}\right\}\Phi_{ml}.$ (7)
Here $L_{\textrm{T}}$ denotes the characteristic locally isothermal length scale set by
@@ 2.2 Disc structure
In order to make things definite, we must assume a particular equilibrium disc structure. Throughout this paper, we will adopt smooth power-law profiles such that the surface density is given by
@@
which yields a temperature profile, $T\propto c_{s}^{2}\propto r^{-q}$. Here and henceforth, all the disc quantities subscripted by ‘p’ are characteristic values evaluated at the planetary semi-major axis. Assuming vertical hydrostatic balance, the aspect ratio of the disc follows
@@
where $H(r)=c_{s}(r)/\Omega_{\textrm{K}}(r)$ is the pressure scale height and $\Omega_{\textrm{K}}(r)=\sqrt{GM_{*}/r^{3}}$ denotes the local Keplerian velocity. Therefore, we can identify $c_{\rm s,p}=h_{\rm p}a_{\rm p}\Omega_{\rm k,p}$.
@@ 2.3 Solving the equations
Having derived the master equation (2.1), encapsulating the excitation of the density wake, we will now briefly summarize our solution method. Whilst purely analytical techniques, such as that pioneered by Goldreich & Tremaine (1979, 1980), typically consider the local wave excitation process in the vicinity of resonance locations, we instead numerically solve for the global structure between $r_{\rm in}=0.05a_{\rm p}$ and $r_{\rm out}=5a_{\rm p}$. This approach was first described by Korycansky & Pollack (1993) and proved successful in subsequent circular planet-disc interaction studies (see Miranda & Rafikov, 2019a, 2020) where the same master equation manifested. More recently, this technique has been extended towards the eccentric planetary regime in FR22 and FR25, where the essential difference is that the Fourier decomposition of the perturbing potential depends on two quantum numbers, $(m,l)$. Therefore, one must compute the wave response for many more forcing modes. Intuitively, $m$ governs the spatial detail of the wake and thus we will require many modes to sufficiently capture the fine structure. Meanwhile, $|m-l|$ measures departures from a circular orbit, for which $m=l$. Hence the more eccentric orbits, and thus more time-dependent wakes, demand higher ‘off-diagonal’ $l$ contributions. In practice, we extract each $\Phi_{ml}$ contribution via a numerical double integral across a grid of radii, using a trapezoidal quadrature with 1024 points in both $\phi$ and $t$. We then interpolate onto arbitrary $r$ as required by the ordinary differential equation (ODE) solver. The master equation is integrated out from corotation resonances, where $\omega_{ml}=\Omega$, and outgoing radiative boundary conditions are applied to fix the contributions from the homogeneous and forced wave components. Additional corrections, such as the phase-gradient error minimisation procedure, slightly adjust the boundary conditions in order to reduce contamination from inwardly propagating wave components, effectively maximising the transmission coefficients of outgoing waves (for details see Korycansky & Pollack, 1993).

### v2

@@ Abstract
Modelling the gravitational interaction between an eccentric perturber and a differentially shearing gas disc is a longstanding problem with various astrophysical applications, ranging from the evolution of planetary systems to the migration of black holes embedded in AGN discs. Recent work has advanced a global, linear, modal approach for calculating the excited wake and the resulting feedback on the perturber’s orbital evolution. In this work we perform a complementary suite of targeted hydrodynamic simulations to test this linear framework across a range of disc temperature and density profiles. In particular, we push from circular orbits to highly eccentric trajectories for which the perturber moves supersonically with respect to the background gas. We find remarkable agreement between our simulations and the linear methodology across a range of diagnostics – lending support to the predicted wake morphologies, complex radial torque density profiles, and torque reversal phenomena, which occur when the eccentricity exceeds the local aspect ratio. In contrast, comparison with previous fitting functions exposes noticeable discrepancies, cautioning against their indiscriminate use in studies which explore a wide range of perturber eccentricities, in varied disc structures. Our simulations also probe the fundamentally nonlinear effects of shock induced angular momentum deposition and coorbital horseshoe drags, which exhibit clear variations with eccentricity. Finally, this careful comparison between linear theory and numerics provides a detailed benchmark for planet-disc interaction problems and therefore we have provided a repository of our linear calculations for use as a rigorous test of future numerical investigations.
@@ 1 Introduction
Whilst the majority of attention has focused on the regime where the perturber is on a circular orbit, there is evidence to suggest that eccentric orbits might also be common. Indeed, observed spiral arm pitch angles in protoplanetary discs are sometimes difficult to explain using a single, circular planet (e.g. Monnier et al., 2019; Uyama et al., 2020). Furthermore, the presence of multiple spiral pattern speeds in SAO 206462 is suggestive of two planets or an eccentric perturber (Xie et al., 2021). These planet-disc interactions might also connect with the post-disc-dispersal exoplanet demographics, which exhibit a clear eccentric population (e.g. Kane et al., 2012; Xie et al., 2016; Eylen et al., 2019; Debras et al., 2021). The origin for these eccentricities might owe to a variety of mechanisms as described in the recent review by Paardekooper et al. (2023), including planet-planet scatterings (e.g. Ford & Rasio, 2008; Lega et al., 2013), secular or resonant gravitational interactions with the surrounding disc (e.g. Goldreich & Sari, 2003; Teyssandier & Ogilvie, 2016; Ragusa et al., 2018) and thermal torques (e.g. Masset, 2017; Eklund & Masset, 2017).
@@
Understanding how such eccentricities modify the wave launching process, and associated torque feedback on the perturber, is therefore important. In this work we will focus our attention on the low mass planetary regime where we can gain rigorous analytical insight by linearizing the fluid equations. The foundational work in this Type-I regime was laid out by Goldreich & Tremaine (1979, 1980). They argued that the wave excitation and coupling to the planetary potential is localised to Lindblad and corotation resonances. The total torque is then found by a summation over these contributions and only depends on the perturbing potential and disc properties at these discrete locations. Subsequent generalizations refined this picture by accounting for gradients in disc properties (Ward, 1986) and by reintroducing non-WKB terms (Artymowicz, 1993). Nonetheless, the aforementioned methods are fundamentally local in character. An alternative approach is to solve the full, linearized wave equation across the global disc without further approximation, as described by Korycansky & Pollack (1993). This method has proven successful in previous circular planet-disc studies, highlighting appreciable differences compared with the isolated resonant picture (see Rafikov & Petrovich, 2012; Petrovich & Rafikov, 2012), whilst also capturing subtle features in global torque density structures (e.g. Miranda & Rafikov, 2019a, b; Cimerman et al., 2024). Furthermore, extended local approaches have shed light on three dimensional effects, whilst also accounting for small orbital eccentricities/inclinations, by means of numerical solutions to modified shearing box wave equations (Tanaka et al., 2002; Tanaka & Ward, 2004; Tanaka & Okada, 2024).
Inspired by these successes, this global methodology was recently extended to eccentric perturber orbits by Fairbairn & Rafikov (2022) (hereafter FR22), which found that the linear theory is capable of capturing the detailed, fine structure seen in the complex wake morphologies simulated by Zhu & Zhang (2022). Fairbairn & Rafikov (2025) (hereafter FR25) subsequently employed this machinery, undertaking an expansive study of the torque excitation and associated back-reaction onto the planetary orbital elements. They comprehensively explored a suite of disc profiles and eccentricities up to $0.12$, finding a stark transition in behaviour when the eccentricity exceeds the local disc aspect ratio. They also improved on previous torque and eccentricity evolution prescriptions found in the literature, which were heuristically motivated and often limited in their range of validity (e.g. Papaloizou & Larwood, 2000; Cresswell & Nelson, 2006; Ida et al., 2020).
@@
These tantalising features demand a more targeted study which can be directly compared with the global, linear method of FR25. Therefore, in this paper we will perform a tailored set of 2D simulations which examine planet-disc interaction across a range of disc profiles whilst pushing towards high eccentricity values of $\leq 0.7$. This significantly extends the previous upper limit of 0.12 explored in FR25 and challenges the capabilities of the analytical framework. With these bespoke simulations we will examine the wake morphology, radial excitation of torque and the net feedback on the planet, thereby verifying the linear theory in unprecedented detail. In section 2 we will review the methodology underpinning the linear results, before we describe the numerical setup in section 3. We will then present our results in section 4 and discuss them in section 5. Finally we will conclude in section 6. Additionally, in appendix A we discuss the convergence properties of our linear framework and numerical simulations, whilst appendix B provides a repository of our data for benchmarking future numerical investigations.
@@ 2.1 Basic approach
We adopt a simple two-dimensional, coplanar and inviscid disc model about a star of mass $M_{\star}$, described in terms of the cylindrical coordinates $(r,\phi)$, similar to that employed by Lin & Papaloizou (1986). The axisymmetric equilibrium disc is specified by the radial profiles of surface density $\Sigma$ and the vertically-integrated pressure $P=c_{\textrm{s}}^{2}\Sigma$, where $c_{\rm s}$ is the isothermal sound speed. The equation for radial force balance then specifies the angular velocity $\Omega$, such that the azimuthal velocity $u_{\phi}=r\Omega$ while the radial velocity $u_{r}=0$.
We now consider the linear perturbing effect of a planet with mass $M_{\rm p}$, which is inserted into the disc. This linear approach formally requires that $M_{\rm p}\ll M_{\rm th}$, where the thermal mass $M_{\rm th}=h_{\rm p}^{3}M_{\star}$ and $h_{\rm p}$ denotes the disc aspect ratio at the planetary semi-major axis location. Specifically we set this low mass planet on a fixed Keplerian orbit with eccentricity $e$, semi-major axis $a_{\rm p}$ and resulting mean motion $n_{\rm p}=\sqrt{GM_{\star}/a_{\rm p}^{3}}$. [^2] A schematic of this setup is visualised in Fig.1 of FR22. Throughout this work, we consider the planet-disc interaction in the astrocentric frame centred on the stellar primary and only include the direct perturbing potential,
$\Phi(r,\phi,t)=-\frac{GM_{\text{p}}}{\sqrt{r^{2}+R(t)^{2}-2rR(t)\cos\left[\phi-\psi(t)\right]+\epsilon^{2}}}$ (1)
Here, $R(t)$ and $\psi(t)$ denote the time-dependent planetary coordinates whilst $\epsilon$ is a smoothing length, chosen to soften the singularity at the planet location according to a simple Plummer prescription. In this work, we have neglected the effect of the indirect term arising from the non-inertial frame. Whilst including this term can lead to physically interesting and previously unappreciated features in the disc angular momentum flux (e.g. Rafikov et al. in prep, Crida et al., 2025), the dominant wake structure and torque feedback on the planet is not strongly affected (see FR25). Since the orbit is non-circular, one cannot remove the fundamental time-dependence of the direct forcing by simply moving into a frame corotating with the planetary mean motion. Instead, the perturbing potential must be expanded in terms of two Fourier numbers – in azimuthal space and orbital time – such that,
$\Phi=\sum_{m=1}^{\infty}\sum_{l=-\infty}^{\infty}\Phi_{ml}(r)\cos\left(m\phi-ln_{\rm p}t\right).$ (2)
Here, the cosine dependence expresses the freedom to choose the phase of the orbit such that pericentre passage occurs at $t=0$ and the perturbing components $\Phi_{ml}$ are real. This external potential forces a linear response in the fluid variables. The net fluid quantities are therefore written as $X+\delta X$ where $X$ represents the equilibrium background whilst $\delta X$ is the small perturbation. These perturbations are similarly assumed to have a Fourier form,
@@
where $\omega_{ml}=(l/m)n_{\rm p}$ denotes the modal pattern speed (generally different from the planetary mean motion when $l\neq m$). Upon linearizing the governing continuity and Euler equations, then inserting the modal ansatz, one arrives at
@@
where $\kappa^{2}=(2\Omega/r)d(r^{2}\Omega)/dr$ is the squared epicylic frequency and $\tilde{\omega}=m(\omega_{ml}-\Omega)$ is the Doppler shifted forcing frequency. Also note that here we have dropped the subscripts $ml$ for ease of notation. Equations (4)-(6) are closed by specifying an equation of state (EoS). Throughout this paper we adopt the locally isothermal approximation for which the temperature $T_{\rm iso}(r)\propto c_{s}^{2}$ is a fixed function of radius and $\delta P=c_{s}^{2}\delta\Sigma$. [^3] The variable $T$ will represent a number of quantities throughout this paper (e.g. temperatures, torques and timescales), so particular care should be taken to note the distinguishing subscript notation. We choose to focus on this simple thermodynamic model so as to directly interface with the comparable approach adopted in FR22 and FR25. However, it should be remembered that alternative EoS prescriptions can lead to different wave behaviour – as seen for adiabatic and ‘beta-cooling’ discs by Miranda & Rafikov (2020). Finally, combining equations (4)-(6) in favour of the pseudo-enthalpy variable, $\delta h=\delta P/\Sigma$, eventually yields the master, modal, linear equation,
$\frac{d^{2}}{dr^{2}}\delta h_{ml}$ $+\left\{\frac{d}{dr}\ln\left(\frac{r\Sigma}{D}\right)-\frac{1}{L_{T}}\right\}\frac{d}{dr}\delta h_{ml}$
@@
$\left.+\frac{1}{L_{T}}\frac{d}{dr}\ln\left(\frac{r\Sigma}{L_{T}D}\right)+\frac{m^{2}}{r^{2}}+\frac{D}{c_{\textrm{s}^{2}}}\right\}\delta h_{ml}=$
@@
$+\left\{\frac{2m\Omega}{r\tilde{\omega}}\left[\frac{d}{dr}\ln\left(\frac{\Sigma\Omega}{D}\right)\right]+\frac{m^{2}}{r^{2}}\right\}\Phi_{ml},$ (7)
which should be solved for each combination of $(m,l)$. Here $L_{\textrm{T}}$ denotes the characteristic locally isothermal length scale set by
@@ 2.2 Disc structure
In order to make things definite, we must assume a particular equilibrium, unperturbed disc structure. Throughout this paper, we will adopt smooth power-law profiles such that the surface density is given by
@@
which yields a temperature profile, $T_{\rm iso}\propto c_{s}^{2}\propto r^{-q}$. Here and henceforth, all the disc quantities subscripted by ‘p’ are characteristic values evaluated at the planetary semi-major axis. Assuming vertical hydrostatic balance, the aspect ratio of the disc follows
@@
where $H(r)=c_{s}(r)/\Omega_{\textrm{K}}(r)$ is the pressure scale height and $\Omega_{\textrm{K}}(r)=\sqrt{GM_{*}/r^{3}}$ denotes the local Keplerian velocity. Therefore, we can identify $c_{\rm s,p}=h_{\rm p}a_{\rm p}n_{\rm p}$.
@@ 2.3 Solving the equations
Having derived the master equation (2.1), encapsulating the excitation of the density wake, we will now briefly summarize our solution method. Whilst purely analytical techniques, such as that pioneered by Goldreich & Tremaine (1979, 1980), typically consider the local wave excitation process in the vicinity of resonance locations, we instead numerically solve for the global structure between $r_{\rm in}=0.05a_{\rm p}$ and $r_{\rm out}=5a_{\rm p}$. This approach was first described by Korycansky & Pollack (1993) and proved successful in subsequent circular planet-disc interaction studies (see Miranda & Rafikov, 2019a, 2020) where the same master equation manifested. More recently, this technique has been extended towards the eccentric planetary regime in FR22 and FR25, where the essential difference is that the Fourier decomposition of the perturbing potential depends on two quantum numbers, $(m,l)$. Therefore, one must compute the wave response for many more forcing modes. Intuitively, $m$ governs the spatial detail of the wake and thus we will require many modes to sufficiently capture the fine structure. Meanwhile, |ml||m-l| measures departures from a circular orbit, for which $m=l$. Hence the more eccentric orbits, and thus more time-dependent wakes, demand higher ‘off-diagonal’ $l$ contributions. In practice, we extract each $\Phi_{ml}$ contribution via a numerical double integral across a grid of radii, using a trapezoidal quadrature with 1024 points in both $\phi$ and $t$. We then interpolate onto arbitrary $r$ as required by the ordinary differential equation (ODE) solver. The master equation is integrated out from corotation resonances, where $\omega_{ml}=\Omega$, and outgoing radiative boundary conditions are applied to fix the contributions from the homogeneous and forced wave components. Additional corrections, such as the phase-gradient error minimisation procedure, slightly adjust the boundary conditions in order to reduce contamination from inwardly propagating wave components, effectively maximising the transmission coefficients of outgoing waves (for details see Korycansky & Pollack, 1993).